Two years ago the Krusty Krab Restaurant purchased a grill for $50,000. The owner, Eugene Krabs, has learned that a new grill is available that will cook Krabby Patties twice as fast as the existing grill. This new grill can be purchased for $100,000 and would be depreciated straight line over 8 years, after which it would have no salvage value. Eugene Krab expects that the new grill will produce EBITDA of $50,000 per year for the next eight years while the existing grill produces EBITDA of only $35,000 per year. The current grill is being depreciated straight line over its useful life of 10 years after which it will have no salvage value. All other operating expenses are identical for both grills. The existing grill can be sold to another restaurant now for $30,000. The Krusty Krab's tax rate is 40%. If in the following tasks a number should be stated with negative sign, enter a minus sign "-" before the number. Don't use a comma as a thousand separator for the results. a. The incremental free cash flow that the Krusty Krab will incur today (Year 0) if they elect to upgrade to the new grill is $ . (round to the nearest dollar) b. The incremental free cash flow that the Krusty Krab will incur in year 1 if they elect to upgrade to the new grill is $ . (round to the nearest dollar) c. If the Krusty Krab's opportunity cost of capital is 6%, then the NPV for upgrading to the new grill is $ . (round to the nearest dollar) d. The IRR for upgrading to the new grill is %. (round to one decimal)
Answer #1

a. Cash flow in year 0 = -100000 (purchase cost of new grilll) + 30000 (sale of od equipment)+0.4(40000 -30000) [tax wite off od grill sd at loss] , 40000 is book value after dep for 2 years,i.e,50000/10*2years = 10000 is dep for old grill

CF in year 0 =-100000+30000+4000 = - $66000.

b. Incremental EBITDA = 50000-35000 =15000

Incremental depreciation = 100000/8 -50000/10 = 7500

Incremental cash flow = (Inc EBITDA-Inc Dep)(1-tax rate)+ Inc Dep

= (15000-7500)(1-0.4) + 7500

= $ 12000

c. From (a) and (b)

Incremental cash flow at year 0 =66000

Increamental cash flows in year 1 = 12000 , period = 8years, discount factor (cost of capita) = 6%

NPV = -66000+ PVIFA (12000 , 8y,6%)

= -66000+12000(6.21)......., PV FACTOR OF 6% FOR 8 YEARS = 6.21

= -66000+74520

= $8520

D. At IRR , Investmwnt - pv of cash fows = 0.i.e, NPV =0

Investmwent = 66000

Cash flows for year = 12000

period = 8years

rate = ?

Try discounting at 18%,  NPV = -66000+ PVIFA (12000 , 8y,8%) = -66000+12000(5.75) = 3000

Try discounting at 18%,  NPV = -66000+ PVIFA (12000 , 8y,10%) = -66000+12000(5.33) = -2040

for 2% decrease , value increased to 5040

for ?% , value increased to 2040(to make it to 0)   

= 2*2040/5040 = 0.81

IRR = 10-0.81 =9.19%